# Template-Based Visual Program Distillation

### v2

@@ 4 Experiments
For all experiments, the teacher model is GPT-4o-mini, the student model is CodeT5 Wang et al. (2023) and is trained with LoRA Hu et al. (2021). We use the GQA dataset Hudson and Manning (2019) and VQAv2 Goyal et al. (2017) datsets for our experiments. We evaluate on the full GQA test-dev split and randomly sample 10000 questions from the VQAv2 validation split. GQA is evaluated using exact match and VQAv2 is evaluated based on annotator/answer agreement as in the original benchmark.
We use a slightly modified API from the original ViperGPT paper  Surís et al. (2023). The main differences are some additional functions (to reduce program length) and removal of the use of a VLM if earlier parts of a program fail. Please see Appendix Section C for more details. For visual models, we use InstructBLIP (Flan-T5 XL) Dai et al. (2023) for general visual queries, Owl-ViT2 Minderer et al. (2024) for detection and CLIP Radford et al. (2021) for classification.
Auto-Context Generation
Before training the student model, we adapt the teacher to the VQA domain. In Table 4, we investigate the effectiveness of using human-generated programs and programs generated by the teacher model validated by answer correctness. As our aim is to enable visual programming with low cost and effort, we compare using a small number (25) of human-generated programs to generating 1000 programs (yielding hundreds of validated programs).
Dataset Human Generated Auto-Context Generated Performance
@@
0 286 (out of 500) 60.3
Table 4: Teacher performance compared across varying numbers of human-generated and auto-context-generated in-context examples.
@@
Data Usage GQA VQAv2
Aug. No Aug. Aug. No Aug
0.1% 41.9 (+0.8) 41.1 61.1 (+0.3) 60.8
0.05% 39.7 (+4.1) 35.6 59.9 (+9.1) 50.8
0.02% 34.2 (+4.2) 30.0 49.0 (+7.8) 41.2
Table 5: Answer accuracy with and without augmentation across different dataset sizes: 1000 (0.1%), 500 (0.05%), and 250 (0.02%) for GQA and 500 (0.1%), 250 (0.05%) and 125 (0.02%) for VQAv2. Performance gain from augmentation is higher when using 0.05% and 0.02% of the dataset compared to 0.1%.
Augmentation vs Non-Augmentation
When training the student model, we use templates and variable substitution to create a more diverse set of programs for training. We evaluate the effect of template-based data augmentation in two ways. In Table 5, we analyze answer accuracy as a function of varying dataset sizes. The results indicate that data-augmentation increases the answer accuracy of the student model, with the largest benefits when few generated programs are available for training.
To better evaluate the effectiveness of distillation, we also evaluate whether the teacher and student provide the same answer. In Table 6, we see that augmentation leads to much higher rates in student-teacher agreement, even when the overall improvement in accuracy is small (see Table 5).
GQA VQA
With Augmentation 64.5 76.1
No Augmentation 59.8 72.0
Table 6: Student/Teacher Prediction Agreement. Data augmentation has a much larger effect on student/teacher prediction compared to answer accuracy.
Cost and Efficiency
One of our objectives is keeping annotation costs low. As shown in Table 7, the total annotation cost for program auto-generation on both datasets is less than a dollar. GQA costs a bit more because of the larger size and during augmentation, GPT-4o-Mini is called when arguments do not appear in the question. By using a small model, inference speed also greatly increases.
GQA VQAv2
Annotation Cost \$0.69 \$0.26
CodeT5 Inference Time (q/s) 35 32.7
Llama-3.2-1B Inference Time (q/s) 0.32 0.33
GPT-4o Mini Inference Time (q/s) 1.3 1.3
Table 7: Annotation cost and inference time for several different models. Our method has low cost and high inference speed with CodeT5.
Comparison with Related Work and Program Evaluation
We compare our approach with the most closely related setting: prompt-based visual program generation. Other works, such as VisRep Khan et al. (2024), have different objectives and use carefully curated and larger datasets, and thus are not included in our comparisons. Furthermore, VisRep evaluates on selected subsets of various datasets, and the specific subsets, models, and code have not been made publicly available for comparison.
Method GQA VQA
Answer Acc Program Acc Answer Acc Program Acc
GPT-4o-Mini (w/Human Examples) 38.5 78 54.2 92
Llama 3.2 1B (w/Human Examples) 19.0 22 23.3 40
Code-T5 Distilled w/o Augmentation (Ours) 41.1 48 60.8 79
Code-T5 Distilled w/Augmentation (Ours) 41.9 68 61.1 80
GPT-4o-Mini (Auto-context) (Ours) 43.1 85 60.3 93
Table 8: Answer and program accuracy of different methods on GQA and VQAv2. With augmentation, distilled programs can achieve similar or better answer accuracy as prompting methods and improve program accuracy compared to not using augmentation.
The widely followed ViperGPT Surís et al. (2023) approach is to use a prompt containing an API and a set of human generated in-context examples. We use two models to evaluate prompting GPT-4o-Mini and Llama-3.2-1B Dubey et al. (2024), a distilled version of Llama 3. 25 human generated in-context examples are used per dataset for each model. In Table 8, we refer to these baselines as GPT-4o-Mini (w/human examples) and Llama-3.2-1B (w/human examples). From our method, we evaluate on two distilled CodeT5 models with and without augmentation. The distilled models are trained with 0.1% of the data (474 question/answer pairs for GQA and 286 question/answer pairs for VQAv2). We also include the auto-context teacher model, notated as GPT-4o-Mini (Auto-context).
Table 9: Common errors and examples used during qualitative evaluation.
Error Question Program Explanation
Not executable Is the chair on the left or right side? chair = image_patch.find(‘chair’) side = choose_relationship(chair, image_patch, ‘left or right’) choose_relationship requires a list as input
API Violation What color is the running dog on the left? left = crop_position(‘running left’, image_patch) dog = left.find(‘dog’) color = dog.classify(‘color’) ‘Running left’ is not a direction or preposition
Contradicts Question What color is the car above the road? road = image_patch.find(‘road’) below_road = crop_position(‘below’, road) car = below_road.find(‘car’) car_color = car.classify(‘color’) The question states that the car is above the road, not below.
Does Not Answer Question Are there two tables? tables = image_patch.find(‘table’) num_tables = count(tables) The question asks if there are two tables, not how many tables there are.
Does not include all question information Is the blue toy small? toy = image_patch.find(‘toy’) is_small = toy.verify_property(‘small’) answer = is_small The question specifically asks about the blue toy.
We evaluate using two metrics: answer accuracy and program accuracy. Program accuracy involves a human (the authors) manually evaluating each program (without execution) for correctness. We randomly sample 100 questions from each dataset for evaluation. The same 100 questions are used to evaluate each method.
@@
  • Not Executable: The program must be executable and return the correct data type (a string for VQA datasets).
  • API Violation: Visual programming APIs are designed to follow basic visual knowledge and reasoning. For example, the ‘find’ function is used with nouns while ‘verify_property’ is generally used for attributes. Clear violations such as trying to find an attribute (e.g. ‘find(green)’) or cropping with a verb (e.g. ‘crop_position(running)’) are considered incorrect.
  • Contradicts Question: Programs that assume a statement that directly conflicts with a statement in the question. For example, assuming an object is on the left, when the question states it is on the right.
@@
We can draw several observations from the results in Table 8. From the answer accuracy results, we see that with a small amount of data and a strong enough teacher, student models can achieve similar performance to the auto-context trained teacher. Auto-context generation also improves the teacher performance. The performance from prompting Llama is particularly low, consistent with previous findings showing that it is difficult to transfer prompts across LLMs Sclar et al. (2024) of similar size, let alone much smaller ones.
The program accuracy results are bit more surprising. For all of the methods, program accuracy is higher than answer accuracy especially for GPT-4o-Mini (Auto-context) which has 85% vs 43% and 93% vs 60% accuracy for GQA and VQAv2 (respectively). Data augmentation has a larger effect on program accuracy than question accuracy but a substantial gap still remains between the student and teacher models, indicating more room for improvement. Across the two datasets, we see a much higher overall program accuracy for VQAv2 than for GQA. Both datasets have fairly easy questions but VQAv2 has many that are correctly answered by either calling a vision-language model (the simple_query) function or simple counting, which involves two functions (find and count). Taken together, our evaluation suggests that the quality of the visual programs is much better than what answer accuracy alone would imply. We include additional analysis of program errors in Appendix A.
@@ 5 Conclusion
Our experiments show with template-based visual program distillation, only a small number of answer annotations are needed to achieve similar performance to prompt-based methods. Auto-context generation removes the burden of human generated program annotations while still retaining the same performance. Template-based augmentation improves both student/teacher prediction agreement and program accuracy compared to non-augmented approaches. The distilled student models achieve similar answer accuracy as the teacher at a fraction of the time and cost.
In conclusion, we show that the visual program ability of LLMs can be distilled into much smaller models with only a few question/answer pairs and no human-generated programs. For less than \$6, (\$5 for student training on the cloud and \$1 for annotation), a 770-M coding model can become a visual program generator. We anticipate that use of template-based visual program distillation will enable users and researchers to iterate more quickly on various visual programming systems and broaden their applications.
6 Limitations
There are several limitations to our work:
Indirect Evaluation Answer accuracy is an indirect evaluation of program correctness. As indicated by our program accuracy results shown in Section 4, answer-based evaluation is only a loose indicator of visual program correctness. Promising areas for future work include automatic program accuracy evaluation and closing the gap between answer accuracy and program accuracy.
Reliance on teacher model and API Our method relies on the quality of the teacher model, which in turn is dependent on the quality of the API and general prompt. A future challenge is how to learn from a weaker teacher and/or unreliable API.
Teacher Data Efficiency Following prior works, if a program was incorrect, it was not used. However, LLMs have the ability to self-correct given the appropriate feedback. In settings where question/answer pairs are limited, such an approach could be more cost-effective than discarding the examples. Areas for future work include incorporating self-correction methods or using program correction models such as VDebugger Wu et al. (2024).
Program Execution Time Our method decreases program generation time, but not program execution time. With a single GPU, program execution time on average was 3.37 s and varies quite widely. The long execution time remains a bottleneck in different applications of visual programming.
Limited Program Complexity Existing VQA datasets are fairly simple and most work on visual programming is limited to tasks where the programs can be generated in a single step by an LLM. Most real world applications are multi-step and would require more complex reasoning and knowledge skills than in the evaluation datasets.
@@ Acknowledgement
Appendix A Program Analysis
@@
In Figure 3, we show the frequency of different error types for each method on the GQA dataset. Note that incorrect programs can fall into multiple categories but in this classification, each incorrect program was counted only once. There are very few execution errors and most errors come from either contradicting the question or not answering it. Many of the Llama generated programs returned yes/no even if the question asked for a different string. Augmentation largely reduces such errors but also generates new errors.
@@
In Figure 4, we show 9 generated programs from 3 questions in the GQA dataset and 3 models: student model trained without augmentation, with augmentation and the auto-context teacher. Generated programs for existence questions about a single object and a single attribute like the one in the first column are almost always correct, even for the worse performing model, Llama-3.2-1B. In the second column, we see an example where the model trained without augmentation leaves out details mentioned in the question but the augmented model generates the correct program. All of the programs are incorrect in the last column but for different reasons. Both of the student models use the same program (apart from variable names) and make two mistakes. First the question asks about ‘not warm’ instead of ‘warm’ and the second is that the answer should be an object, not yes/no. The teacher program returns an object but still fails to recognize that the object should be ‘not warm’ even though the variable name includes ‘not warm’ in it. Questions involving negative properties are almost always missed by the teacher and student models.
@@
Appendix B Training and Model Details

### v4

@@ 4 Experiments
For all experiments, the teacher model is GPT-4o-mini. Several student models are trained with LoRA Hu et al. (2022): Qwen2.5-Coder-0.5B Hui et al. (2024), CodeT5 Wang et al. (2023) and Llama-3.2-1B Dubey et al. (2024). We use the GQA dataset Hudson and Manning (2019) and VQAv2 Goyal et al. (2017) datasets for our experiments. All student models are trained and evaluated on a single dataset unless indicated by the word ‘joint’. We evaluate on the full GQA test-dev split and randomly sample 10,000 questions from the VQAv2 validation split. GQA is evaluated using exact match and VQAv2 is evaluated based on annotator/answer agreement as in the original benchmark. Experiments on additional datasets can be found in Appendix A.
Dataset Human Generated LLM Generated Performance
@@
Table 4: Teacher performance compared across varying numbers of human-generated and LLM generated in-context examples (i.e. auto-context generation). Auto-Context Generation achieves the same or better performance than human-generated program annotations.
Error Example Explanation
Not executable Q: Is the chair left or right?
P: side = choose_relationship (chair, img, [‘left’,‘right’])
choose_relationship requires a list as input
API Violation Q: What color is the running dog on the left?
P: crop_position (‘running left’,img)
“running left” isn’t a valid direction/preposition.
Contradicts Question Q: What color is the car above the road?
P: below_road = crop_position(‘below’, road)
The question states that the car is above the road, not below.
Does Not Answer Question Q: Are there two tables?
P: num = count(image.find(‘table’))
The question asks if there are two tables, not how many tables there are.
Does Not Include All Question Information Q: Is the blue toy small?
P: toy.verify_property(‘small’)
The question specifically asks about the blue toy.
Table 5: Common errors encountered during qualitative evaluation.
We use a slightly modified API from the original ViperGPT paper  Surís et al. (2023). The main differences are some additional functions (to reduce program length) and removal of the use of a VLM if earlier parts of a program fail. Please see Appendix D for more details. For visual models, we use InstructBLIP (Flan-T5 XL) Dai et al. (2023) for general visual queries, Owl-ViT2 Minderer et al. (2024) for detection and CLIP Radford et al. (2021) for classification.
Method GQA VQA
Answer Acc Prog. Acc Stu/Tea Agree. Answer Acc Prog. Acc Stu/Tea Agree.
Few-shot prompting
Llama 3.2-1B 19.0 22 23.3 40
Qwen Coder 2.5-0.5B 18.1 14 25.0 16
Teacher
GPT-4o-Mini 43.1 85 60.3 93
Distilled Students w/Aug.
Qwen Coder 2.5-0.5B (single) 42.5 71 78.8 60.8 73 74.2
Qwen Coder 2.5-0.5B (joint) 43.1 78 78.2 60.3 79 70.5
Code-T5-770M (single) 41.9 68 64.5 61.1 80 76.1
Llama 3.2-1B (single) 43.1 73 81.2 60.3 82 73.6
Table 6: Comparison of few-shot prompting and student models trained w/ augmented distillation across different metrics. Few-shot open, source models have poor performance but template-augmented student distillation has comparable performance to the teacher model, GPT-4o-Mini.
Auto-Context Generation
Before training the student model, we adapt the teacher to the VQA domain. We investigate the effectiveness of using human-generated programs and programs generated by the teacher model validated by answer correctness. As our aim is to enable visual programming with low cost and effort, we compare using a small number (25) of human-generated programs to generating 1,000 programs (yielding hundreds of validated programs).
@@
Evaluation Metrics
All approaches are evaluated with two metrics: answer accuracy and program accuracy. Student models are also evaluated on student/teacher answer agreement or if student and teacher programs return the same answer, regardless of correctness. Program accuracy involves a human (the authors) manually evaluating each program (without execution) for correctness. We randomly sample 100 questions from each dataset for program correctness evaluation. The same 100 questions are used to evaluate each method. To ensure the fairness of our program evaluation, we enlist a second annotator to evaluate program correctness on 50 examples: 25 correct and 25 incorrect. The second annotator agreed with the original program evaluation on 47 out of 50 of the examples. Program annotations have been released so others can contribute. [^1] https://github.com/michalsr/vp_distill_templates/tree/main/program_correctness_annotations
@@
- Not Executable The program must be executable and return the correct data type (a string for VQA datasets).
- API Violation Visual programming APIs are designed to follow basic visual knowledge and reasoning. For example, the ‘find’ function is used with nouns while ‘verify_property’ is generally used for attributes. Clear violations such as trying to find an attribute (e.g. ‘find(green)’) or cropping with a verb (e.g. ‘crop_position(running)’) are considered incorrect.
- Contradicts Question Programs that assume a statement that directly conflicts with a statement in the question. For example, assuming an object is on the left, when the question states it is on the right.
@@
Few-Shot Comparison
First, we compare our distilled student models with the most closely related setting: prompt-based visual program generation. Other works, such as VisRep Khan et al. (2024), have different objectives and use carefully curated and larger datasets, and thus are not included in our comparisons. Furthermore, VisRep evaluates on selected subsets of various datasets, and the specific subsets, models, and code have not been made publicly available for comparison.
Two models, Llama-3.2-1B and Qwen2.5-Coder-0.5B, are evaluated using both few-shot prompting and template-based distillation. CodeT5 is also evaluated with template-based distillation. There are two versions of distilled Qwen2.5-Coder-0.5B: one version is trained (and evaluated) on GQA and VQA separately (denoted by ‘single’) and the other is trained on both datasets (denoted by ‘joint’).
The same 25 human generated in-context examples (incorporated into the ViperGPT Surís et al. (2023) API) used in the Auto-Context Generation experiments above are used per dataset for each few-shot model. The distilled models are trained with 0.1% of the data (474 question/answer pairs for GQA and 286 question/answer pairs for VQAv2). All program annotations come from the teacher model.
Method GQA VQA
Answer Acc Prog. Acc Stu/Tea Agree. Answer Acc Prog. Acc Stu/Tea Agree.
Qwen Coder 2.5-0.5B w/o Aug (single) 41.8 69 73.0 60.2 75 72.2
Qwen Coder 2.5-0.5B w/ Aug (single) (+0.7)42.5 (+2)71 (+5.8)78.8 (+0.6)60.8 (-2)73 (+2.0)74.2
Qwen Coder 2.5-0.5B w/o Aug (joint) 41.7 69 73.5 58.8 72 67.6
Qwen Coder 2.5-0.5B w/ Aug (joint) (+1.4) 43.1 (+9)78 (+4.7)78.2 (+1.5)60.3 (+6)78 (+2.9)70.5
Code-T5-770M w/o Aug (single) 41.1 48 59.8 60.8 79 72.0
Code-T5-770M w/ Aug (single) (+0.8)41.9 (+20)68 (+4.7)64.5 (+0.3)61.1 (+1)80 (+4.1)76.1
Llama 3.2-1B w/o Aug (single) 40.0 61 70.1 60.2 79 72.6
Llama 3.2-1B w/ Aug (single) (+3.1)43.1 (+12)73 (+11.1)81.2 (+0.1)60.3 (+3)82 (+1)73.6
Table 7: Effect of augmentation. On average, template-based augmentation improves performance on nearly all models, particularly on student/teacher agreement and program accuracy. Bold results are statistically significant.
Few-Shot Results We can draw several observations from the results in Table 6. Comparing different methods, we see that few-shot prompting on small open-source models results in extremely low performance but with a small amount of data and a strong enough teacher, the same models achieve similar performance to the auto-context teacher, especially on answer accuracy. Program accuracy improves significantly from distillation but a gap still remains indicating room for improvement.
The results across different metrics are a bit more surprising. On all of the student models, the average difference between student/teacher answer agreement and answer accuracy is 33% for GQA and 13% for VQAv2, indicating that answer accuracy underestimates distillation performance. The noisiness of answer accuracy is illustrated even more by program accuracy performance.
The program accuracy results are the most surprising. For all of the methods, program accuracy is higher than answer accuracy especially for the teacher model GPT-4o-Mini which has 85% vs. 43% and 93% vs. 60% program vs answer accuracy for GQA and VQAv2 (respectively). The significant difference among the metrics indicates that most errors in visual programming systems with proprietary LLMs are not from the programs but from the API and visual models. An additional analysis of program errors is in Appendix B.
Data Augmentation
Next, we ablate the effect of data augmentation on distillation. For each model, we train with and without distillation on both datasets and measure performance across the three metrics. From the results in Table 7, we see that data augmentation improves performance across all metrics for nearly all models and datasets. Augmentation has more of an effect on GQA compared to VQAv2. Both datasets have fairly easy questions, but VQAv2 has many questions that are correctly answered by either calling a vision-language model (the simple_query) function or simple counting, which involves two functions (find and count). The relative increase is more notable for program accuracy (average 6.4) and student/teacher agreement (average 4.5) compared to answer accuracy (average 1.1), again indicating that answer accuracy does not fully capture model behavior.
Cost and Efficiency
One of our objectives is low annotation costs and fast inference time. As shown in Table 8, the total annotation cost for program auto-generation on both datasets is less than a dollar. GQA costs a bit more because of the larger size and during augmentation, GPT-4o-Mini is called when arguments do not appear in the question. By using a small model, inference speed also greatly increases.
Metric GQA VQAv2
Teacher
Program Generation Cost (Total) \$0.69 \$0.26
GPT-4o-Mini Inference Time (q/s) 1.3 1.3
Distilled Students
Qwen2.5-Coder Inference Time (q/s) 39.2 40.4
CodeT5 Inference Time (q/s) 35.0 32.7
Llama-3.2-1B Inference Time (q/s) 16.1 18.6
Table 8: Annotation cost and inference time for student and teacher models. Student inference time is much faster than teacher inference time.
@@ 5 Conclusion
Our experiments demonstrate models trained with template-based visual program distillation can become specialized and efficient visual program generators at a small cost. Auto-context generation removes the burden of human generated program annotations while still retaining the same performance. The results also show how commonly used metrics for visual programming, do not fully capture the performance. Human program verification reveals that on the best models, programs are likely not the source of errors and that future work should focus on the API and visual models, not program generation. For less than \$6, (\$5 for student training on the cloud and \$1 for annotation), a 500-M coding model can become a visual program generator. We anticipate that the use of template-based visual program distillation will enable users and researchers to iterate more quickly on various visual programming systems and broaden their use for targeted applications.
6 Limitations and Future Work
There are several limitations and areas for future work:
Specialization. Our method distills a specialized VQA model that is not intended to provide the same breadth of capability as the original LLM. Our specialized models, however, are much faster and cheaper for inference and easy to produce, making them suitable for targeted application.
Effort Required for Program Accuracy. While program accuracy is an important metric, it also requires a significant amount of human effort, making it difficult to evaluate on a large scale. A promising area for future work includes automatic evaluation of program accuracy.
Reliance on teacher model and API. Our method relies on the quality of the teacher model, which in turn is dependent on the quality of the API and general prompt. A future challenge is how to learn from a weaker teacher and/or unreliable API.
Teacher Data Efficiency. Following prior works, if the answer returned by a program was incorrect, it was not used. However, LLMs have the ability to self-correct given the appropriate feedback. In settings where question/answer pairs are limited, such an approach could be more cost-effective than discarding the examples. Areas for future work include incorporating self-correction methods or using program correction models such as VDebugger Wu et al. (2024).
Program Execution Time. We evaluate the time to generate programs, but the time to execute with API calls to several visual models, can be much greater (3.4s / query in our implementation with high variance) and requires significant engineering effort to make efficient.
Limited Program Complexity. Existing VQA datasets are relatively simple and most work on visual programming is limited to tasks where the programs can be generated in a single step by an LLM. Most real world applications are multi-step and would require more complex reasoning and knowledge skills than in the evaluation datasets.
7 Societal Implications
Distilled student models inherit existing biases from LLMs so care is required when deploying the models in production. There are also privacy risks, so safeguards should be taken to prevent data leakage.
@@ Acknowledgement
Appendix A Additional Results
Additional experimental results on two vision-centric datasets, the 2D-split of CV Bench Tong et al. (2024) and the VizWiz Gurari et al. (2018) validation set, can be found in Table 9. We evaluate an augmented student model (Qwen2.5-Coder-0.5B) and the teacher model. For CV Bench Tong et al. (2024), the student model is trained on both GQA and VQAv2 (referred to as ‘joint’ in Section 4). 2000 examples ( 0.1%) of the VizWiz training set are annotated by the teacher model and used to train the student. Like the experiments on VQAv2, we use BERT Devlin et al. (2019) and BART Lewis et al. (2020a), for augmentation. The results show the efficacy of our method on datasets related to the training set as well as real-world tasks.
Method CV Bench VizWiz
Answer Acc Stu/Tea Agree. Answer Acc Stu/Tea Agree.
GPT-4o-Mini (Teacher) 56.0 N/A 34.3 N/A
Qwen Coder 2.5-0.5B (w/ Aug) 54.5 93.0 34.9 71.6
Table 9: Additional evaluation of augmented distillation on the 2D split of CV Bench and the validation split of the real-world dataset VizWiz. The distilled augmented model has similar performance as the teacher model.
N-Gram Entropy
In our experiments, we primarily use task-level metrics to validate our method and assess augmented data quality. To directly quantify data diversity, we also measure N-Gram Entropy Zhang et al. (2018) on 20,000 samples from both our augmented and non-augmented GQA training sets. The augmented data has an entropy of 7.10 compared to 6.18 for the original set, confirming that our augmentation procedure meaningfully increases linguistic variety.
Probability of Augmentation
In Section 3.2, we state that during augmentation, each argument has a 50% probability of being replaced. We evaluate using different percentages on a small validation set of GQA and found 50% had the highest performance, 43.2%, compared to 42.5% for 90% replacement rate and 42.4% for a 10% replacement rate.
Model Size
The focus of our work is to design a method for a small amount of training data as well small models that can be run on consumer GPUs, which led us to focus on models with no more than 1B parameters. However, there are models that are slightly larger, such as 1.5B which leads to the question, whether there are any performance gains from using such models. On a subset of GQA, we evaluate different sizes of Qwen2.5-Coder on 0.5B, 1.5B and 3B. The results, shown in Table 10, show our augmentation method is actually more effective on smaller models. Scaling our method for larger open source models is an area for future work.
Method GQA Subset
Answer Acc Stu/Tea Agree.
Qwen Coder 2.5-0.5B (w/Aug) 43.2 70.5
Qwen Coder 2.5-1.5B(w/ Aug) 42.5 65.8
Qwen Coder 2.5-3B(w/ Aug) 42.4 67.6
Table 10: When using $0.1\%$ of the training dataset, our method is more effective on smaller models,
Appendix B Program Analysis
@@
In Figure 4, we show the frequency of different error types for Qwen-Coder-0.5B under few-shot model, non-augmented (joint) model and augmented (joint) model as well as the teacher model on the GQA dataset. Note that incorrect programs can fall into multiple categories but in this classification, each incorrect program was counted only once. Other than the few-shot model, there are no execution errors. The main sources of error are not answering question and API violation.
@@
In Figure 5, we show 9 generated programs from 3 questions in the GQA dataset and 3 models: CodeT5 student model trained without augmentation, CodeT5 with augmentation and the auto-context teacher. Generated programs for existence questions about a single object and a single attribute like the one in the first column are almost always correct, even for few-shot methods. In the second column, we see an example where the model trained without augmentation leaves out details mentioned in the question but the augmented model generates the correct program. All of the programs are incorrect in the last column but for different reasons. Both of the student models use the same program (apart from variable names) and make two mistakes. First the question asks about ‘not warm’ instead of ‘warm’ and the second is that the answer should be an object, not yes/no. The teacher program returns an object but still fails to recognize that the object should be ‘not warm’ even though the variable name includes ‘not warm’ in it. Questions involving negative properties are almost always missed by the teacher and student models.
@@
Appendix C Training and Model Details